• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Dezern/Fansler

    Log entry time 14:15:09 on April 25, 2007

    Entry number 200006

    keyword=Magnet Lead Flows

    All magnet lead flows began to drop. We tried to ramp the magnets down before the levels reached the trip points. The Q3's and Q1's on both arms tripped before we could reach zero amps. We contacted Cryo, they advised that our return pressure was high, and they would bring a third compressor online to lower our pressures. After the compressor came online, we reset the Q1's from the counting house, and entered Hall A in controlled access to reset both arm Q3's. All magnets are operational at this time. Lead flows are stable.

    A copy of this log entry has been emailed to: fansler@jlab.org, stevensm@jlab.org